OpenStudy (anonymous):

solve for x x^4/5 - 3 =1

OpenStudy (anonymous):

What part are you stuck on?

OpenStudy (anonymous):

i dont understand at all :/

OpenStudy (anonymous):

First take all constants to one side and express the equation in the form of: x^4 = ?

OpenStudy (anonymous):

im sorry i dont get it

OpenStudy (anonymous):

You need to take one step at a time: x^4/5 - 3 = 1 Add 3 to both sides. You get: x^4/5 = 4 Multiply by 5 on both sides. You get: x^4 = 20 Now, you need to deal with exponents. When you take an exponent to the other side, it gets reversed. So, you get: x = 20^(1/4)